# LLMs, You Can Evaluate It! Design of Multi-perspective Evaluation for Reports in Security Operation Centers

### v2

@@ 4.1 Design Motivation and System Requirements
Our motivation for MESSALA is to provide cognition and domain knowledge of veteran SOC practitioners with LLMs. The Analyst-wise Checklist is utilized to realize this idea.
However, it is still insufficient because LLM-based judgments of texts often differ from evaluations of human experts [murugadoss2025evaluating]. Indeed, when we conducted post-interviews with five of the fifteen participants in the interviews with respect to the evaluation of analysis reports by an LLM using the Analyst-wise Checklist, we obtained similar comments as ineffective as the existing work [murugadoss2025evaluating] (“The feedback remains largely superficial and does not appear to provide a sufficiently in-depth evaluation of the report.”, “The comments avoid direct wording, which makes it difficult to translate the feedback into concrete actions.”). Namely, to provide appropriate evaluation results, an LLM must judge a report consistently with evaluations of veteran SOC practitioners and synthesize these judgments into feedback [zhou2024llm]. To do this, we formulate two elements, i.e., numerical scores for quantitative evaluation and feedback for qualitative evaluation, for the evaluation of analysis reports, which correspond to RQ2 and RQ3. MESSALA is also the answers to these questions.
@@ 4.2 Key Idea
The key idea of MESSALA is to imitate SOC practitioners’ cognition [zhong2016automate] for evaluations of analysis reports because it enables LLMs to generate outputs similar to humans [saha2025learning, chu2025think]. To realize this idea, we introduce two new methods, Granularization Guideline and Multi-perspective Evaluation LLM, in addition to the Analyst-wise Checklist. MESSALA can then generate evaluation results and feedback identical to those of veteran SOC practitioners for analysis reports. We describe each method below.
The Granularization Guideline steers the LLM’s inference with expertise and practical feedback of veteran SOC practitioners by translating the Analyst-wise Checklist into actionable evaluations. This enables an LLM to focus on the specific parts, that require detailed evaluations of veteran SOC practitioners, in analysis reports.
@@ 4.3.1 Granularization Guideline
The Granularization Guideline is designed to granularize items to be evaluated in the Analyst-wise Checklist for a given analysis report, and their underlying evaluation criteria. Its design follows prior works on SOC practices [kersten2023give, kerstenfield2025tier1analyst, turcotte2025automated, zhong2016automate]. Based on this design, categories in the guideline were iteratively refined through evaluations of real-world analysis reports and their criteria defined in Section 3, and then were finalized by confirmation with two SOC practitioners from the authors’ affiliations.
@@
Table 7: Overview of selected categories in the Granularization Guideline.
@@
Figure 2: Granularization guideline. The Granularization Guideline can break down items in the checklist into specific evaluation steps for each category.
@@ 4.3.2 Multi-perspective Evaluation LLM with High-level and In-depth Evaluations
Figure 4: Overview of Our Prompt for Granularization and Evaluation. The red text indicates the distinctive characteristics of each prompt.
@@
High-level Evaluation LLM: This allows an LLM to rely on simple evaluation items with superficial information. The correlation between LLM outputs and human ratings for such an evaluation is high [murugadoss2025evaluating], while they often produce holistic evaluations due to the lack of detailed guidelines [zhou2024llm]. Consequently, the evaluation results of LLMs may differ from those of human SOC practitioners. The prompt used for high-level evaluation is shown in Fig. 4. It consists of three components, i.e., the target analysis report, specific items in the Analyst-wise Checklist, and the evaluation setup, including the task and its method. The evaluation is performed using a five-point Likert scale, which can provide human-like analysis [murugadoss2025evaluating] as well as the reliability of a rating scheme [kusmaryono2022number]. We note that, within the prompt, only a minimal evaluation method is provided by following the existing prompt structures [fu2023gptscore].
In-depth Evaluation LLM: This enables an LLM to understand important descriptions. This means that the LLM evaluates analysis reports by approximating the domain-specific contexts as SOC practitioners using the Granularization Guideline. The In-depth Evaluation LLM offers several advantages. First, by interpreting the evaluation process correctly, the LLM can extract rich context from the prompt and easily justify its output. Second, it can automatically granularize each item using LLMs [liu2023g-eval, liu2024hd]. Nevertheless, we note that evaluations using general LLMs may differ from human judgment due to domain-specific contexts [chiang2023closer] and prompts whose detailed instructions are sometimes ineffective [murugadoss2025evaluating]. To avoid such a situation, the prompt used for the In-depth Evaluation LLM consists of three components: an analysis report to be evaluated, the Analyst-wise Checklist, and the Granularization Guideline, as shown in Fig. 3,  4. The task of this LLM is to granularize each item of the checklist in accordance with the given analysis report.
Integration: This step integrates the high-level and in-depth evaluations to evaluate analysis reports. Since each evaluation has distinct characteristics, integrating its outputs yields more reliable and interpretable results. It can also mitigate potential blind spots and biased judgments arising from a single viewpoint by virtue of cross-referencing outputs from different levels of evaluation[liu2024hd]. The prompt is structured as follows. First, the LLM is instructed to generate an initial score for each item in the checklist granularized by the In-depth Evaluation LLM. Next, the LLM is given the results of the High-level Evaluation LLM, and then reflects both results to reconsider the initial scores. The LLM then outputs a final score with a five-point Likert scale and feedback for the given report. The prompt example is shown in Fig. 4.
@@ 5.1.1 Datasets
Concretely, we provided the LLM with (i) attack patterns from MITRE ATT&CK, (ii) typical patterns of detection, conclusion, and response extracted from past internal reports, and (iii) knowledge about hypothetical critical assets. In addition, for each target attack technique, we had the LLM refer to external technical reports so that the generated analyses and countermeasures would better reflect environment- and situation-specific interpretations. These steps were implemented by combining prompt chaining with a RAG-style retrieval process. After generating a large number of pseudo reports through this pipeline, we performed automatic quality scoring using an LLM and iteratively discarded low-quality samples. Finally, we conducted a manual quality review and selected 10 pseudo reports, which we used as the second dataset in our evaluation experiments. Samples of the generation prompts and pseudo-reports are provided in Fig. 5.
@@
Using the Analyst-wise Checklist in Section 3, we were also able to obtain the evaluation results from the five participant in the user study in Section 3. Each item in the Analyst-wise Checklist was rated on a five-point Likert scale, and the final score was computed as the average of their individual scores. To align our understanding of the evaluation criteria, the first twenty reports were jointly reviewed. The remaining thirty reports were then processed by the two authors. For this human gold evaluation, 10 to 15 checklist items were selected, focusing on important contexts of the analysis reports. This limitation in the number of the items is based on two reasons: 1) Evaluating too many items would increase the workload on the participants and may cause low accuracy; 2) In realistic operational settings, it is more practical to concentrate on key items aligned with the content of the analysis reports in order to provide meaningful feedback.
In constructing the expert evaluation dataset, we followed three important precautions: 1) To minimize bias, evaluators were assigned only to reports from SOC environments they were not directly responsible for. As a result, each report was typically scored by approximately three evaluators, and their average was used as the final expert rating. 2) To ensure consistency among evaluators, we conducted preliminary briefings on the checklist items and held a Q&A session using Microsoft Teams to align understanding. 3) To preserve independence during the evaluation process, evaluators were not allowed to view each other scores until all evaluations were completed.
@@ 5.2 Results
We present the results below, comparing the numerical scores of MESSALA with the baseline methods to answer RQ2. An overview of the results is shown in Table 8 and Fig. 7. According to the table, except for a few cases, i.e., RMSE on gpt-4o and qwen3 (14B), $\tau$ on gpt-4.1 mini, and $r$ on gemma3 (12B), MESSALA outperforms the baseline methods across all settings and consistently achieves high evaluation scores. These results highlight the advantages obtained from our two key modules, i.e., 1) the Analyst-wise Checklist and the Granularization Guidelines, and 2) the Multi-perspective Evaluation LLM. Likewise, Fig. 7 shows that MESSALA exhibits the distribution closest to the human gold evaluation. These results demonstrate that MESSALA is significantly effective for evaluations of analysis reports because it is more correlated with human judgment than the other methods.
@@
Figure 6: LLM-generated feedback examples per method with sanitized reports and checklist items.
@@ 6.1.1 Experimental Setting
We limit the length of feedback comments to 500 tokens by summarizing the results derived from each method. In this study, since multiple checklist items are evaluated for a single report, presenting all evaluation results as-is would result in an excessive amount of information. Accordingly, we chose to follow the appropriate amount of feedback indicated in prior studies [zhou2024llm, liang2024can].
The summarized feedback comments are created as critical comments, focusing on points that received particularly low evaluation scores. Although best practices [hattie2007power] for feedback generally recommend including positive comments to support user motivation, analysis reports often require rapid actions. Therefore, feedback comments in this evaluation are restricted to checklist items that are less than or equal to “3" with the following prompt: “Based on the following evaluation results, generate feedback comments for improving the report in no more than 500 tokens. Preserve as much information from the original evaluation as possible, describe the issues concretely, and do not add any new information.”
@@ 6.2.1 Defect-injected Analysis Reports
The defect-injected analysis report dataset is designed to evaluate whether an LLM can accurately identify defects in analysis reports. It consists of reports with realistic defects that are observed in SOC environments: for instance, missing contextual information, insufficient justification for judgments, flawed causal reasoning, and unclear or impractical response descriptions. The dataset also defines defect categories with their reference review comments. These reference review comments are used to assess whether the LLM can generate appropriate feedback.
We define a defect as at least one of the eleven evaluation perspectives derived in Section 3 is missing or unsatisfied. Based on defect categories in prior works [liu2023reviewergpt, d2024marg] and the reference review comments collected from SOCs of the authors’ organization, we consolidated the original eleven evaluation perspectives into four defect categories with confirmations by SOC practitioners. Table 11 shows the consolidation between the four defect categories and the eleven perspectives.
@@
In addition, following these categories, we inject defects into the reports, and it is desirable for the LLM to generate feedback that captures the reference review comments corresponding to the defects. Such reference review comments are derived from real analysis reports, for example: “The report does not explain why the communication occurred, nor does it assess whether the communication or the involved endpoint is anomalous.” We selected a subset of analysis reports of high quality from standpoints of both human analysts and an LLM to inject defects. Each report was then modified according to predefined rules to inject one or more defects from the defect categories described above. During the above injections, only the relevant parts of each analysis report were modified, while the overall structure and writing style were preserved.
These datasets were constructed using LLMs, whereby their outputs were reviewed and refined by the authors and some defect samples were further confirmed through reviews by SOC practitioners. We finally utilized 44 analysis reports as the dataset. Furthermore, we conduct a preliminary validation to verify that quality deterioration of analysis reports caused by the injected defects is measurable. For each defect category, we compare evaluation scores between clean and defect-injected analysis reports derived from the same original report, and then identify whether score deteriorates in the corresponding evaluation perspectives while scores on unrelated evaluation perspectives remain unchanged. Defect injection and validation are conducted independently and the LLM performs scoring without access to defect labels or human verification results. The results are presented in Table 12.
Each defect category does not correspond to a single checklist item in a one-to-one manner. Instead, a defect typically affects multiple quality dimensions simultaneously, such as justification, contextualization, and decision support. Therefore, we intentionally avoid defining strict mappings between defect categories and checklist items.
While the defects are artificially injected, they are intentionally designed to be subtle and context-dependent, rather than exaggerated or synthetic. Moreover, multiple defects may coexist within a single report, reflecting the ambiguous and compound nature of quality issues in real SOC analysis reports. This design choice ensures that the dataset captures the inherent uncertainty and complexity of real-world report evaluation.
Finally, this dataset focuses on evaluating the ability to identify and assess report defects, which is a prerequisite for generating meaningful improvement suggestions. An example of a defect-injected report is shown in Fig. 9.
@@ 6.2.3 Results
Method 1 evaluates the report as sufficient for the checklist item related to alert causation, primarily because it mentions the roles of the devices and their associated vulnerabilities.
In contrast, MESSALA assigns low evaluation scores when the rationale supporting the conclusion is insufficient. Based on granularized evaluation items, MESSALA identifies and explicitly points out that the report fails to explain how the detected device anomalies or state changes would impact building operations.
@@
Method 1 is able to partially identify specific defects in the report; however, its feedback is limited to high-level observations and lacks practical guidance. It does not clearly indicate which analytical evidence is missing or how analysts should re-examine the underlying assumptions, resulting in feedback that is insufficient for improving report quality.
Compared with Method 1, MESSALA provides more informative and practical feedback by explicitly identifying insufficient analytical reasoning and inadequately justified false-positive determinations. Rather than accepting high-level operational explanations at face value, MESSALA critically examines the relationship between detected anomalous communication patterns and legitimate business traffic, thereby revealing detailed analytical gaps that Method 1 fails to capture.

### v3

@@ 4.1 Design Motivation and System Requirements
Our motivation for MESSALA is to incorporate the cognition and domain knowledge of veteran SOC practitioners into LLM-based evaluation. The Analyst-wise Checklist is utilized to realize this idea. However, using the checklist alone is still insufficient because LLM-based judgments of texts often differ from evaluations of human experts [murugadoss2025evaluating]. Indeed, when we conducted post-interviews with five of the fifteen participants in the interviews with respect to the evaluation of analysis reports by an LLM using the Analyst-wise Checklist, we obtained feedback similar to that reported in prior work [murugadoss2025evaluating] , such as:“The feedback remains largely superficial and does not appear to provide a sufficiently in-depth evaluation of the report.”, “The comments avoid direct wording, which makes it difficult to translate the feedback into concrete actions.”
Therefore, to provide appropriate evaluation results, an LLM must judge a report consistently with evaluations of veteran SOC practitioners and synthesize these judgments into feedback [zhou2024llm]. To do this, we formulate two elements, i.e., numerical scores for quantitative evaluation and feedback for qualitative evaluation, for the evaluation of analysis reports, which correspond to RQ2 and RQ3. MESSALA is designed to address these questions.
@@ 4.2 Key Idea
The key idea of MESSALA is to imitate the cognitive process used by SOC practitioners [zhong2016automate] when evaluating analysis reports, as this enables LLMs to generate outputs similar to humans [saha2025learning, chu2025think]. To realize this idea, we introduce two new methods, Granularization Guideline and Multi-perspective Evaluation LLM, in addition to the Analyst-wise Checklist. MESSALA can then generate evaluation results and feedback identical to those of veteran SOC practitioners for analysis reports. We describe each method below.
The Granularization Guideline steers the LLM’s inference with expertise and practical feedback of veteran SOC practitioners by translating the Analyst-wise Checklist into actionable evaluations. This enables an LLM to focus on the specific parts that require detailed evaluations of veteran SOC practitioners, in analysis reports.
@@ 4.3.1 Granularization Guideline
The Granularization Guideline is designed to granularize items to be evaluated in the Analyst-wise Checklist for a given analysis report, and their underlying evaluation criteria. Its design follows prior works on SOC practices [kersten2023give, kerstenfield2025tier1analyst, turcotte2025automated, zhong2016automate]. Based on this design, categories in the guideline were iteratively refined through evaluations of real-world analysis reports using the Analyst-wise Checklist items in Section 3, and then were finalized by confirmation with two SOC practitioners from the authors’ affiliations.
@@
Table 7: Overview of categories defined in the Granularization Guideline.
@@
Figure 2: Illustration of the Granularization Guideline in use. The guideline illustrates how checklist items are broken down into concrete evaluation steps for each category.
@@ 4.3.2 Multi-perspective Evaluation LLM with High-level and In-depth Evaluations
Figure 4: Overview of Prompts for Granularization and Evaluation. The red text indicates the distinctive characteristics of each prompt.
@@
High-level Evaluation LLM: This allows an LLM to rely on simple evaluation items with superficial information. The correlation between LLM outputs and human ratings for such an evaluation is high [murugadoss2025evaluating], while they often produce holistic evaluations due to the lack of detailed guidelines [zhou2024llm]. Consequently, the evaluation results of LLMs may differ from those of human SOC practitioners. The prompt used for high-level evaluation is shown in Fig. 4. It consists of three components, i.e., the target analysis report, specific items in the Analyst-wise Checklist, and the evaluation setup, including the task and its method. The evaluation is performed using a five-point Likert scale, which can provide human-like judgments [murugadoss2025evaluating] as well as the reliability of a rating scheme [kusmaryono2022number]. We note that, within the prompt, only a minimal evaluation method is provided by following the existing prompt structures [fu2023gptscore].
In-depth Evaluation LLM: This enables an LLM to understand important descriptions. This means that the LLM evaluates analysis reports by approximating the domain-specific contexts as SOC practitioners using the Granularization Guideline. The In-depth Evaluation LLM offers several advantages. First, by interpreting the evaluation process correctly, the LLM can extract rich context from the prompt and easily justify its output. Second, it can automatically granularize each item using LLMs [liu2023g-eval, liu2024hd]. Nevertheless, we note that evaluations using general LLMs may differ from human judgment due to domain-specific contexts [chiang2023closer] and prompts whose detailed instructions are sometimes ineffective [murugadoss2025evaluating].
To avoid such a situation, the prompt used for the In-depth Evaluation LLM consists of three components: an analysis report to be evaluated, the Analyst-wise Checklist, and the Granularization Guideline, as shown in Fig. 3 and  4. The task of this LLM is to granularize each item of the checklist in accordance with the given analysis report.
Integration: This step integrates the high-level and in-depth evaluations to evaluate analysis reports. Since each evaluation has distinct characteristics, integrating their outputs yields more reliable and interpretable results. It can also mitigate potential blind spots and biased judgments arising from a single viewpoint by virtue of cross-referencing outputs from different levels of evaluation[liu2024hd]. The prompt is structured as follows. First, the LLM is instructed to generate an initial score for each item in the checklist granularized by the In-depth Evaluation LLM. Next, the LLM is given the results of the High-level Evaluation LLM, and then reflects both results to reconsider the initial scores. The LLM then outputs a final score with a five-point Likert scale and feedback for the given report. The prompt example is shown in Fig. 4.
@@ 5.1.1 Datasets
Concretely, we provided the LLM with (i) attack patterns from MITRE ATT&CK, (ii) typical patterns of detection and response extracted from past analysis reports, and (iii) knowledge about hypothetical critical assets. In addition, for each target attack technique, we had the LLM refer to external reports so that the generated analysis reports would reflect interpretations specific to environments and situations. These steps were implemented by combining prompt chaining with a RAG-style retrieval process. After generating a large number of pseudo-reports through this pipeline, we used an LLM to automatically score quality and iteratively discarded low-quality samples. Finally, we conducted a manual quality review and selected 10 pseudo-reports, which we used as the second dataset in our evaluation experiments. Samples of the generation prompts and pseudo-reports are provided in Fig. 5.
@@
Using the Analyst-wise Checklist in Section 3, we were also able to obtain the evaluation results from the five participants in the user study in Section 3. Each item in the Analyst-wise Checklist was rated on a five-point Likert scale, and the final score was computed as the average of their individual scores. To align our understanding of the evaluation criteria, the first twenty reports were jointly reviewed. The remaining thirty reports were then processed by the two authors. For this human gold evaluation, 10 to 15 checklist items were selected, focusing on important contexts of the analysis reports. This limitation in the number of items is based on two reasons: 1) Evaluating too many items would increase the workload on the participants and may cause low accuracy; 2) In realistic operational settings, it is more practical to concentrate on key items aligned with the content of the analysis reports in order to provide meaningful feedback.
In constructing the expert evaluation dataset, we followed three important precautions: 1) To minimize bias, evaluators were assigned only to reports from SOC environments they were not directly responsible for. As a result, each report was typically scored by approximately three evaluators, and their average was used as the final expert rating. 2) To ensure consistency among evaluators, we conducted preliminary briefings on the checklist items and held a Q&A session using Microsoft Teams to align understanding. 3) To preserve independence during the evaluation process, participants were not allowed to view each other scores until all evaluations were completed.
@@ 5.2 Results
We present the results below, comparing the numerical scores of MESSALA with the baseline methods to answer RQ2.
An overview of the results is shown in Table 8 and Fig. 7. According to the table, except for a few cases, i.e., RMSE on gpt-4o and qwen3 (14B), $\tau$ on gpt-4.1 mini, and $r$ on gemma3 (12B), MESSALA outperforms the baseline methods across all settings and consistently achieves high evaluation scores. These results highlight the advantages obtained from our two key modules, i.e., 1) the Analyst-wise Checklist and the Granularization Guidelines, and 2) the Multi-perspective Evaluation LLM. Likewise, Fig. 7 shows that MESSALA exhibits the distribution closest to the human gold evaluation. These results demonstrate that MESSALA is significantly effective for evaluations of analysis reports because it is more correlated with human judgment than the other methods.
@@
Figure 6: LM-Generated Evaluation Examples for Each Method Using Sanitized Reports and Checklist Items.
@@ 6.1.1 Experimental Setting
We limit the length of feedback comments to 500 tokens by summarizing the results derived from each method. In this study, since multiple checklist items are evaluated for a single report, presenting all evaluation results as-is would result in an excessive amount of information. Accordingly, we chose to follow the appropriate amount of feedback indicated in prior studies [zhou2024llm, liang2024can]. The summarized feedback comments are created as critical comments, focusing on points that received particularly low evaluation scores. Although best practices [hattie2007power] for feedback generally recommend including positive comments to support user motivation, analysis reports often require rapid actions. Therefore, feedback comments in this evaluation are restricted to checklist items that are less than or equal to “3". The following instructions are given in the prompt: “Based on the following evaluation results, generate feedback comments for improving the report in no more than 500 tokens. Preserve as much information from the original evaluation as possible, describe the issues concretely, and do not add any new information.”
@@ 6.2.1 Defect-injected Analysis Reports
The defect-injected analysis report dataset is designed to evaluate whether an LLM can accurately identify quality defects in SOC analysis reports. The dataset consists of reports with realistic, context-dependent defects commonly observed in SOC environments, such as missing contextual information, insufficient justification for judgments, flawed causal reasoning, and unclear or impractical response descriptions. Formally, we define a defect as a case where at least one of the eleven evaluation perspectives derived in Section 3 is missing or unsatisfied.
Based on defect categories proposed in prior work [liu2023reviewergpt, d2024marg] and reference review comments collected from SOCs within the authors’ organization, we consolidated the original eleven evaluation perspectives into four defect categories, with confirmation by SOC practitioners. Table 11 summarizes the relationship between the four defect categories and the original evaluation perspectives.
@@
Following these categories, we inject one or more defects into high-quality analysis reports and define reference review comments that explicitly point out the inserted defects. These reference review comments are derived from real SOC analysis reports, for example: “The report does not explain why the communication occurred, nor does it assess whether the communication or the involved endpoint is anomalous.” To construct defect-injected reports, we first selected a subset of analysis reports that were evaluated as high quality by both human analysts and an LLM. Each report was then modified according to predefined injection rules, while preserving the overall structure and writing style by only altering the relevant portions.
These datasets were constructed with the assistance of LLMs, whose outputs were subsequently reviewed and refined by the authors, and some defect samples were further validated through reviews by SOC practitioners. We finally utilized 44 analysis reports as the dataset. Furthermore, we conduct a preliminary validation to verify that quality deterioration of analysis reports caused by the injected defects is measurable. For each defect category, we compare evaluation scores between clean and defect-injected analysis reports derived from the same original report, and then identify whether score deteriorates in the corresponding evaluation perspectives while scores on unrelated evaluation perspectives remain unchanged. Defect injection and validation are conducted independently and the LLM performs scoring without access to defect labels or human verification results. The results are presented in Table 12.
Importantly, defect categories do not correspond to checklist items in a one-to-one manner. A single defect often affects multiple quality dimensions simultaneously, such as justification, contextualization, and decision support. Therefore, we intentionally avoid defining strict mappings between defect categories and checklist items. Although the defects are artificially injected, they are designed to be subtle and compound, reflecting the ambiguous and intertwined nature of quality issues in real SOC analysis reports. This dataset focuses on evaluating the ability to identify and assess such defects, which is a prerequisite for generating meaningful improvement suggestions. An example of a defect-injected report is shown in Fig. 9.
@@ 6.2.3 Results
Method 1 evaluates the report as sufficient for the checklist item related to alert causation, primarily because it mentions the roles of the devices and their associated vulnerabilities. In contrast, MESSALA assigns low evaluation scores when the rationale supporting the conclusion is insufficient. Based on granularized evaluation items, MESSALA identifies and explicitly points out that the report fails to explain how the detected device anomalies or state changes would impact building operations.
@@
Method 1 is able to partially identify specific defects in the report; however, its feedback is limited to high-level observations and lacks practical guidance. It does not clearly indicate which analytical evidence is missing or how analysts should re-examine the underlying assumptions, resulting in feedback that is insufficient for improving report quality. Compared with Method 1, MESSALA provides more informative and practical feedback by explicitly identifying insufficient analytical reasoning and inadequately justified false-positive determinations. Rather than accepting high-level operational explanations at face value, MESSALA critically examines the relationship between detected anomalous communication patterns and legitimate business traffic, thereby revealing detailed analytical gaps that Method 1 fails to capture.